hello i need some help on this thankyou

Mrs. Maaza is 3 times as old as her son. In 12 years Mrs. MAaza will be twice as old. How old is her son now.??

Answer by ilana(307)

Let's say Mrs. Maaza is x years old and her son is y years old.
So x=3y and x+12=2(y+12). Now we have a system of equations to solve.
So plug 3y in for x in the second equation. 3y+12=2y+24 (using distribution).
Solving for y gives y=12. So the son is 12 years old. Let's check. If he is 12, she is 3*12=36. In 12 years, she will then be 48 and he will be 24, so it works:)